OpenStudy (anonymous):

Logarithms. Solve.. Log7 (x-4)-log7 (x+5)=log7 (3) Help please The 7 is the supposed to be the base of log

OpenStudy (zzr0ck3r):

\(\log_7(x-4)-\log_7(x+5)=\log_7(3)\\\log_7(\dfrac{x-4}{x+5})=\log_7(3) \\ \dfrac{x-4}{x+5}=3\)

OpenStudy (anonymous):

Thanks